OpenStudy (anonymous):

Which of the following is not a valid probability distribution for a discrete random variable? Check all that apply. A. 1/5,1/10 ,1/10 ,1/10 ,1/5 ,1/10 ,1/10 , 1/10 B. 1/3,1/4 ,1/5 ,1/6 C. 1/2,1/4 ,1/8 ,1/16 ,1/32 ,1/64 ,1/128 ,1/128 D. -1/2, -1/3, -1/4, -1/5, 137/60 E. 1/6, 1/6, 1/6, 1/6, 1/6, 1/6

OpenStudy (tkhunny):

Discard D. immediately. Negative numbers are not allowed. What other criterion is definitive?

OpenStudy (anonymous):

well anything that you could count because we are looking for measurements

OpenStudy (tkhunny):

?? I don't know what that means. The must add to unity. E. 1/6 six times is 1. Perfect. Try the others.

OpenStudy (anonymous):

well A. is 7/10ths and I'm not to sure about C.

OpenStudy (tkhunny):

A. is not 7/10. The are 6 at 1/10 and 2 at 1/5 = 2/10. That one looks good.

OpenStudy (tkhunny):

C. Is fun. Think backwards on it. We have to get to 1 Start with 1 and subtract things off. 1 - 1/2 = 1/2 1/2 - 1/4 = 1/4 1/4 - 1/8 = 1/8 etc...

OpenStudy (anonymous):

so 1/2, 1/4, 1/8, 1/16, 1/32, 1/64?

OpenStudy (tkhunny):

1/8 - 1/16 = 1/16 <== That's how much there is left. 1/16 - 1/32 = 1/32 <== Now, there is just this much left. 1/32 - 1/64 = 1/64 1/64 - 1/128 = 1/128 1/128 - 1/128 = 0 and we are done. That one summed to unity.

OpenStudy (anonymous):

I think I'm starting to get it now

OpenStudy (anonymous):

How about B. ?

OpenStudy (tkhunny):

You really just have to find a way to add them up. I used decimals. 1/3,1/4 ,1/5 ,1/6 Using hundreths 1/3 is 33ish 1/4 is 25, giving now 58ish 1/5 is 20, giving now 78ish Well, we have 22 hundreths to go. Will 1/6 cover it?

OpenStudy (anonymous):

no it only gives 16 right?

OpenStudy (tkhunny):

Yup. that one is a little short. We used a different method to add each of them. Stay flexible!!

OpenStudy (anonymous):

so what method did you use?

OpenStudy (tkhunny):

All available. Whatever seemed most appropriate for the given set of data.

OpenStudy (anonymous):

alright so A, C, and E are not a valid probability distribution for a discrete random variable?

OpenStudy (anonymous):

Right?

OpenStudy (tkhunny):

Backwards. We disqualified B. and D.

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

oh I get it

OpenStudy (tkhunny):

Are they all positive? Do they sum to unity (1)? Done.